OpenStudy (briellshaw):

If McKayla had 4 table sets each table set came with 6 chairs if McKayla got 13 more table sets how many chairs and tables would she have in all???

OpenStudy (thebadyboy):

well 6 chairs and 1 table 4 sets 13+4= 17 * 7=?

OpenStudy (anonymous):

she has 102 chairs and 17 tables

OpenStudy (briellshaw):

Can you draw me an example

OpenStudy (briellshaw):

@bellahasnochill or @thebadyboy